OpenStudy (anonymous):

will award medal, find the solution to the differential equaiton. 3(du/dt)=u^2 where u(0)=5

OpenStudy (freckles):

have trouble separating the variables?

OpenStudy (freckles):

try dividing u^2 on both sides and multiply dt on both sides

OpenStudy (anonymous):

\[\int\limits1/u^2du=\int\limits1/3dt\]

OpenStudy (anonymous):

is this how you would set up your integration

OpenStudy (freckles):

\[3 du =u^2 dt \\ 3 \frac{1}{u^2} du=dt \\ \text{ or } \frac{1}{u^2} du=\frac{1}{3} dt \] 2nd line or 3rd line works so in other words what you have is good

OpenStudy (anonymous):

so to solve for the equation you have |u^2|=(e^(t/3))e^c where (e^c=b)

OpenStudy (freckles):

where did all e^ come from?

OpenStudy (freckles):

you don't have 1/u du on the left

OpenStudy (anonymous):

when you integrate you have ln|u^2|, and to get rid of the ln you need to raise both sides under e

OpenStudy (freckles):

\[\int\limits_{}^{}u^{-2} du= \frac{u^{-2+1}}{-2+1}+C\]

OpenStudy (freckles):

\[\int\limits_{}^{}u^{-n} du=\frac{u^{-n+1}}{-n+1}+C ; n \neq 1 \\ \text{ when n=1 we have } \int\limits_{}^{}u^{-1} du=\ln|u|+C\]

OpenStudy (anonymous):

ohhh right

OpenStudy (anonymous):

=-1/u

OpenStudy (anonymous):

-1/u=t/3+c

OpenStudy (freckles):

yep yep :)

OpenStudy (anonymous):

c=-1/5

OpenStudy (freckles):

u(0)=5 gives -1/5=0/3+c yes c=-1/5 you are right

OpenStudy (anonymous):

so it asks the question in terms of u=...

OpenStudy (freckles):

so the first step I would make is writing the right hand side as one fraction by combining the fractions

OpenStudy (freckles):

then just flip both sides

OpenStudy (anonymous):

u=-(15/5t-3)

OpenStudy (anonymous):

hell yea thanks

OpenStudy (freckles):

np